An empty 1.0 L was filled with 5.0 moles of oxygen and 4.0 moles of nitrogen monoxide. If the amount of NO2 at equilibrium is 1.5 moles, calculate the equilibrium constant, Kc.

 

2NO(g)  + O2(g)          2 NO2 (g)           

 



1
Expert's answer
2022-02-22T20:47:01-0500

"K_c=\frac{[NO_2]^2}{[NO]^2[O_2]}"


"\begin{matrix}
 & 2NO_{g}&+&O_{2(g)}&\rightleftharpoons&2NO_2\\I&4&&5&&0
 \\C&-1.5&&-0.75&&1.5\\E&2.5&&4.25&&1.5
\end{matrix}"



"K_c=\frac{1.5^2}{(4.25)(2.5)^2}=0.084706"